# Index theory for singular Lagrangian systems and Bessel-type differential operators

### v1

@@ Abstract
The aim of the present manuscript is to develop an index theory for singular Lagrangian systems, with a particular focus on the important class of singular operators given by Bessel-type differential operators. The main motivation is to address several challenges posed by singular operators, which appear in a wide range of applications: celestial mechanics (for instance, perturbations in planetary motion), oscillatory systems with time-dependent forcing, electromagnetism (such as wave equations in nonuniform media), and quantum mechanics (notably certain Schroedinger equations with periodic potentials).
@@ 1 Introduction
In its classical form, the Morse Index Theorem asserts that the number of negative eigenvalues of the Hessian of the geodesic action functional at a critical point equals the number of conjugate points, counted with multiplicity, along the corresponding geodesic. Through standard spectral flow techniques, this integer can be expressed as the sum of the dimensions of the kernels of a one-parameter family of regular Sturm–Liouville operators. For singular Sturm–Liouville operators, however, this framework fails, and the usual spectral flow identities no longer apply.
@@ 1.1 Highlights and Main Results
One of the main contributions of this first part of the paper is the explicit construction of a symplectic subspace $(W,\omega|_{W})$ in terms of $\ker L^{*}$, that is, the space of solutions to the Sturm–Liouville equation (see Section 4 for further details), and of a trace map that we now describe.
@@
It is worth noting that the bracket $[f,g](t)$ coincides with the standard symplectic form once $\mathbb{R}^{2n}$ is endowed with the local coordinate system $(P(t)f^{\prime}(t)+Q(t)f(t),f(t))$.
@@
    $\operatorname{\mathcal{T}}(f):=\big(-[f,y_{1}](a^{+}),\ldots,-[f,y_{2k-2n}](a^{+}),[f,z_{1}](b),\ldots,[f,z_{2n}](b)\big).$
    This case corresponds to the limit circle case or regular case in the Weyl classification of the boundary conditions.
@@
    The space $W$ is defined explicitly in terms of a basis of $\ker L^{*}$. Moreover, since the entire construction depends only on the domains of $L^{*}$ and $L$, the definition of $W$ remains valid if $L$ is replaced by another operator $\widetilde{L}$ (not necessarily belonging to the same family) provided that $D(\widetilde{L})=D(L)$.
@@
    the case of regular Sturm–Liouville operators on bounded domains;
@@
    the case of Sturm–Liouville operators on $\mathbb{R}$$^{+}$.
@@
Using the above spectral flow formula, we prove in Section 5 the Morse index theorem for singular Sturm–Liouville operators.
@@
    (H3) The quadratic form on $L^{2}((a,b),{\mathbb{R}}^{n})$ associated with the SL-operator $l$ is bounded from below.
@@
If $l$ is regular at $b$ and Dirichlet boundary conditions are imposed, then
@@
The case of general Lagrangian boundary conditions is treated in detail in Section 6. In particular, we prove that even in this singular framework, the Morse index with general boundary conditions differs from the Dirichlet case by a correction term computed via the triple index of the Cauchy data space and the projections of the domains of $L_{\Lambda}$ and of the Friedrichs extension $L_{F}$. More precisely, we obtain the following result.
Theorem 4.
Let $L$ be a one-sided singular Sturm–Liouville operator as defined in Equation (3.1), and suppose that conditions (H3)(H4) hold. Let $\Lambda\in\operatorname{\mathrm{Lag}}(W,\omega|_{W})$. Then
where $\iota$ denotes the triple index (cf. Appendix A and the references therein).
Theorem 4 provides a rigorous quantification of a phenomenon first observed by Rellich. More precisely, if $T$ is a closed self-adjoint operator bounded from below, and $A$ is a relatively bounded symmetric perturbation with sufficiently small relative bound, then
$S=T+A$
remains self-adjoint and bounded from below (see [Kat80, Chapter 5, Section 4, Remark 4.13]).
However, this property does not necessarily hold for unbounded perturbations. As Kato explains in [Kat80, Chapter 5, Section 4, Remark 4.13], based on Rellich’s example, when a sequence of self-adjoint operators $T_{n}$ converges in the gap topology to a self-adjoint operator bounded from below, the lower bounds of $T_{n}$ may diverge to -$\infty$, and eigenvalues may “disappear” at -$\infty$. Rellich’s scalar counterexample in the context of Sturm–Liouville operators illustrates this effect, although the precise number of disappearing eigenvalues remains an open problem. Section 7 provides a precise quantification of the multiplicity of eigenvalues that escape to -$\infty$.
@@
Then every self-adjoint extension of $L_{R}$ is a Fredholm operator.
@@ 5 The Morse Index Theorem for Singular Sturm-Liouville
    (H3) The quadratic form on $L^{2}((a,b),{\mathbb{R}}^{n})$ associated to the SL-operator $l$ is bounded from below
@@ 5.1 Preliminary Results on Lagrangian Intersection Theory
Since $\bigcap_{a<\sigma<b}D(t_{L_{(a,\sigma),\Lambda}})=\set{0}$, then there exists $c_{0}\in(a,b)$ such that $\mathrm{n_{-}}(L_{c_{0},\Lambda})=0$, concluding the proof.
@@ 9.1 A description of the problem
Summing up all previous arguments we get a new version and proof of the main result proved by authors at [BHPT20].
@@
The proof follows by the above discussion and by invoking Theorem 8.9 and Theorem 5.2. The last statement is a direct consequence of [BS08]. ∎
@@ Appendix C Friedrichs Extensions: Definition and Basic Results
Let $A:D(A)<H\to H$ be a densely defined, symmetric operator on a Hilbert space $H$, satisfying the semiboundedness condition:
@@ D.2 Morse indices of forms and of Friedrich extensions
Let $\alpha=t_{A_{b,\Lambda}}$ and $V=D(\alpha)$. Then $\alpha(u,v)=\langle A_{b}u,v\rangle_{V}$. We assume that $Z<V$ is the negative spectral subspace of $\alpha$, let $V_{n}=D(t_{A_{a^{+}\dfrac{1}{n},\Lambda}})$ and let $K_{n}=V_{n}^{\alpha}$ where
@@
There exists a bounded Fredholm operator $B_{D}:D(t_{L_{D}})\to D(t_{L_{D}})$ such that $t_{L_{D}}(u,v)=(Bu,v)_{W^{1,2}}$ for each $u,v\in D(t_{L_{D}})$.
Proof.
We start observing that the norm on $D(t_{L_{D}})=W^{1,2}_{0}((a,b),{\mathbb{R}}^{n})$ is equivalent to the norm on $W_{0}^{2,2}((a,b),{\mathbb{R}}^{n})$ defined by $((L_{D}+\lambda I)u,u)_{L^{2}}^{1/2}$ for $\lambda$ sufficiently large.
With respect to this norm, the associated operator $B_{\Lambda_{D}}$ is $L_{\Lambda_{D}}(L_{\Lambda_{D}}+\lambda)^{-1}=I-\lambda(L_{\Lambda_{D}}+\lambda)^{-1}$ and it is a Fredholm operator since $(L_{\Lambda_{D}}+\lambda)^{-1}:W_{0}^{1,2}\to W_{0}^{1,2}$ is compact. ∎
Corollary D.20.
There exists a bounded Fredholm operator $B_{\Lambda}:D(t_{L_{\Lambda}})\to D(t_{L_{\Lambda}})$ such that $t_{L_{\Lambda}}(u,v)=(Bu,v)_{W^{1,2}}$ for each $u,v\in D(t_{L_{\Lambda}})$.
Proof.
Since $t_{L_{\Lambda}}$ is bounded, there exists $B_{\Lambda}$ such that $t_{L_{\Lambda}}(u,v)=(B_{\Lambda}u,v)_{W^{1,2}}$. We have
$(B_{\Lambda}u,v)=(B_{\Lambda_{D}}u,v),u,v\in D(t_{L_{\Lambda_{D}}}).$
Let $T:D(t_{L_{\Lambda}})\to D(t_{L_{D}})$ be the orthogonal projection. Then we have $B_{D}=TB_{\Lambda}T$. So, in particular $\operatorname{im}(B_{D})<\operatorname{im}(B_{\Lambda})$. It means that $T(\operatorname{im}B_{\Lambda})$ has finite codimension. Since $\dim\ker T=\dim D(t_{L_{\Lambda}})/D(t_{L_{D}})\leq 2n$, we have
$\operatorname{im}T/(T\operatorname{im}(B_{\Lambda}))\cong(H/\ker T)/(\operatorname{im}B_{\Lambda}+\ker T)/\ker T)\cong H/(\operatorname{im}B_{\Lambda}+\ker T).$
By this, it follows that $\operatorname{codim}(\operatorname{im}B_{\Lambda})\leq\operatorname{codim}(T\operatorname{im}B_{\Lambda})+2n$. Then we can conclude that $B_{\Lambda}$ is a Fredholm operator. ∎
Lemma D.21.

### v2

@@ Abstract
The aim of the present manuscript is to develop an index theory for singular Lagrangian systems, with a particular focus on the important class of singular operators given by Bessel-type differential operators. The main motivation is to address several challenges posed by singular operators, which appear in a wide range of applications: celestial mechanics (for instance, perturbations in planetary motion), oscillatory systems with time-dependent forcing, electromagnetism (such as wave equations in nonuniform media), and quantum mechanics (notably certain Schrodinger equations with periodic potentials).
@@ 1 Introduction
In its classical form, the Morse Index Theorem asserts that the number of negative eigenvalues of the Hessian of the geodesic action functional at a critical point equals the number of conjugate points, counted with multiplicity, along the corresponding geodesic. Through standard spectral flow techniques, this integer can be expressed as the sum of the dimension of the intersections between the graph of the fundamental solution with the Lagrangian subspace corresponding to the boundary condition. For singular Sturm–Liouville operators, however, this classical framework fails, and the usual spectral flow identities no longer apply.
@@ 1.1 Highlights and Main Results
One of the main contributions of this first part of the paper is the explicit construction of a symplectic subspace $(W,\omega|_{W})$ through the $\ker L^{*}$, that is, the space of solutions to the Sturm–Liouville equation (see Section 4 for further details), and of a trace map that we now describe.
@@
    $\operatorname{\mathcal{T}}(f):=\big(-[f,y_{1}](a^{+}),\ldots,-[f,y_{2k-2n}](a^{+}),[f,z_{1}](b),\ldots,[f,z_{2n}](b)\big),$
    where $y_{i}\in\ker L^{*}$ for $i=1,\ldots,2k-2n$. The existence of this subset $\widetilde{y}_{1},\ldots,\widetilde{y}_{2k-2n}$ of functions in $\ker L^{*}$ is proved at Lemma 4.4 for details. We refer to Remark 4.7 where is fully explained how to concretely select such a subset of functions. This case corresponds to the limit circle case or regular case in the Weyl classification of the boundary conditions.
@@
    The space $W$ is defined explicitly by using a basis of $\ker L^{*}$. Moreover, since the entire construction depends only on the domains of $L^{*}$ and $L$, the definition of $W$ remains valid if $L$ is replaced by another operator $\widetilde{L}$ (not necessarily belonging to the same family) provided that $D(\widetilde{L})=D(L)$.
@@
    Regular Sturm–Liouville operators on bounded domains;
@@
    Sturm–Liouville operators on $\mathbb{R}$$^{+}$
  3. 3.
    Bessel-type operators on $(0,1]$ and $1,\infty)$.
@@
Using the above spectral flow formula, we prove in Section 5 the Morse index theorem for singular Sturm–Liouville operators. For the regular case or the heteroclinic and half-clinic cases, we refer the interested reader to [HWY20] , [HPWX20], [BCJLMS18] and [HP17].
@@
    (H3) The quadratic form on $\mathscr{C}^{\infty}_{0}((a,b),{\mathbb{R}}^{n})$ associated with the SL-operator $l$ is bounded from below in $L^{2}$.
@@
Analogously, if $l$ is regular at $b$ and Dirichlet boundary conditions are imposed, then
@@
The case of general Lagrangian boundary conditions is treated in detail in Section 6. In particular, we prove that even in this singular framework, the Morse index with general boundary conditions differs from the Dirichlet case by a correction term computed via the triple index of the Cauchy data space and the projections of the domains of $L_{\Lambda}$ and of the Friedrichs extension $L_{F}$.
Section 7 is devoted to study a rigorous quantification of a phenomenon first observed by Rellich in the case of singular Sturm-Liouville equations. In the abstract terms the stage is the following. If $T$ is a closed self-adjoint operator bounded from below, and $A$ is a relatively bounded symmetric perturbation with sufficiently small relative bound, then $S=T+A$ remains self-adjoint and bounded from below (see [Kat80, Chapter 5, Section 4, Remark 4.13]).
However, this property does not necessarily hold for unbounded perturbations. As Kato explains in [Kat80, Chapter 5, Section 4, Remark 4.13], based on Rellich’s example, when a sequence of self-adjoint operators $T_{n}$ converges in the gap topology to a self-adjoint operator bounded from below, the lower bounds of $T_{n}$ may diverge to -$\infty$, and eigenvalues may “disappear” at -$\infty$. Rellich’s scalar counterexample in the context of Sturm–Liouville operators illustrates this effect, although the precise number of disappearing eigenvalues remains an open problem. Section 7 provides a precise quantification of the multiplicity of eigenvalues that escape to -$\infty$. In the regular case, this phenomenon has been already studied at [HLWZ19, HLWZ22].
@@
If R(t)>1/4R(t)>-1/4, then every self-adjoint extension of $L_{R}$ is a Fredholm operator.
@@
If $q$ is hyperbolic, then the Morse index $\mathrm{n_{-}}(q)$ of $q$ is always finite.
@@ 4.1 The trace map
In the construction above we defined the trace map, but we did not specify how to select the 2k2n2k-2n functions $\{y_{1},\dots,y_{2k-2n}\}\subset\ker L^{*}$ so that
$W=\operatorname{span}\{y_{1},\dots,y_{2k-2n},\,z_{1},\dots,z_{2n}\}.$
We choose them to make $\omega|_{W}$ nondegenerate. It suffices to ensure that the restriction
$\omega\big|_{\operatorname{span}\{\tilde{y}_{1},\dots,\tilde{y}_{2k-2n}\}}$
is nondegenerate, where $\{\tilde{y}_{i}\}$ is an appropriate reordering of a basis of $\ker L^{*}$. To this end, start with any basis $\{y_{1},\dots,y_{k}\}$ of $\ker L^{*}$ and consider the $k\times k$ Gram matrix $G=(G_{ij})$ of the boundary pairing
$G_{ij}\;=\;[y_{i},y_{j}](a),\qquad 1\leq i,j\leq k.$
Reorder the basis so that the principal $(2k-2n)\times(2k-2n)$ block $\big([y_{i},y_{j}](a)\big)_{1\leq i,j\leq 2k-2n}$ is invertible (this is always possible as soon as $\operatorname{rank}G\geq 2k-2n$). Then take
$\{y_{1},\dots,y_{2k-2n}\}$
to be the first 2k2n2k-2n vectors in this reordered basis. By construction, $\omega$ is nondegenerate on $\operatorname{span}\{y_{1},\dots,y_{2k-2n}\}$, and hence $\omega|_{W}$ is nondegenerate as claimed.
Remark 4.8.
@@ 5 The Morse Index Theorem for Singular Sturm-Liouville
    (H3) The quadratic form on $\mathscr{C}^{\infty}_{0}((a,b),{\mathbb{R}}^{n})$ associated to the SL-operator $l$ is bounded from below in $L^{2}$.
@@ 5.1 Preliminary Results on Lagrangian Intersection Theory
Since $\bigcap_{a<\sigma<b}D(t_{L_{(a,\sigma),\Lambda}})=\set{0}$, then by invoking Lemma D.15 and Lemma C.2, there exists $c_{0}\in(a,b)$ such that $\mathrm{n_{-}}(L_{c_{0},\Lambda})=0$, concluding the proof.
@@ 6 Morse Index Theorem for General Boundary Conditions
The difference between two operators subject to different boundary conditions has been investigated in [HWY20] and [HPWX20], both for regular Sturm–Liouville operators and for Sturm–Liouville operators defined on the entire real line. In [BCLS24, Theorem 1.2], this result has been further generalized to arbitrary symmetric operators.
Here, by employing an approach analogous to that developed in [HWY20], we establish a difference formula for general Sturm–Liouville operators subject to distinct boundary conditions.
@@ 9.1 A description of the problem
Summing up all previous arguments we get a new proof of the main result proved by authors at [BHPT20, HOY21].
@@
If $q$ is hyperbolic, then the Morse index $\mathrm{n_{-}}(q)$ of $q$ is always finite.
@@
The proof follows by the above discussion and by invoking Theorem 8.9 and Theorem 5.2. The second statement is a direct consequence of [BS08]. The last statement readily follows from Theorem 8.8. (Cf. [HOY21] for further details). ∎
@@ Appendix C Friedrichs Extensions: Definition and Basic Results
Let $(H,\langle\cdot,\cdot\rangle$ be a real Hilbert space and let $A:D(A)<H\to H$ be a densely defined, symmetric operator on a Hilbert space $H$, satisfying the semiboundedness condition:
@@ D.2 Morse indices of forms and of Friedrich extensions
Let $\alpha=t_{A_{b,\Lambda}}$ and $V=D(\alpha)$. Then $\alpha(u,v)=\langle A_{b}u,v\rangle_{V}$. We assume that $Z<V$ is the negative spectral subspace of $\alpha$, let $V_{n}=D(t_{A_{a+\frac{1}{n},\Lambda}})$ and let $K_{n}=V_{n}^{\alpha}$ where